OpenStudy (anonymous):

The perimeter of a triangle is 60 inches. If the sides of the triangle are 2x, 3x, and 5x, find the length of the shortest side. A.) 1 ft B.) 2 in. C.) 4 in. D.) 6 in.

OpenStudy (anonymous):

@welshfella

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Nope, that's why I put it on here!

OpenStudy (skullpatrol):

Let a = 2x, b = 3x, and c = 5x.

OpenStudy (anonymous):

x=10?

OpenStudy (skullpatrol):

where a, b, and c are the lengths of the sides of the triangle.

OpenStudy (skullpatrol):

a + b + c = 60

OpenStudy (anonymous):

So 6?

OpenStudy (skullpatrol):

yes x = 6 :)

OpenStudy (skullpatrol):

The question asks for "find the length of the shortest side."

OpenStudy (skullpatrol):

a = 2x, b = 3x, and c = 5x. a = 2*6 = 12 in is the shortest side

OpenStudy (skullpatrol):

recall 12 in = 1 ft